OpenStudy (anonymous):

@SolomonZelman Could you see my answer to this question and see if I got it correct?

OpenStudy (anonymous):

ill try to check for u

OpenStudy (anonymous):

Ok here it is.

OpenStudy (anonymous):

1. The main tank has a radius of 70 feet. What is the volume of the quarter-sphered sized tank? Round your answer to the nearest whole number. You must explain your answer using words, and you must show all work and calculations to receive credit. In the description, we are told that we need to find the volume of a quarter-sphered tank, so in order to do that we need to use this formula (Volume of a sphere): V = 4/3 3.14 r3. Now that we know what formula to use, lets apply this to our problem. V = 4/3 (3.14) (r)3. V = 4/3 (3.14) (70)3. V = 4/3 (3.14) (343000). V = 4/3 (1077,020) V = 1,436,026.67 In the description, we are told to find the volume of a “quarter sized sphere”. We have just found the volume for the whole sphere, so in order to find the volume for a quarter of it we need to divide our total volume by four. 1,436,026.67 / 4 = 359,006.6675 Now we round our answer to the nearest whole number and get a grand total of 359,007. Now we know that the volume of the quarter sized sphere tank is 359,007. Lets move on to the calculation!

OpenStudy (anonymous):

Just read past the question marks, there is nothing under them.

OpenStudy (anonymous):

im not sure im not good at word problems

OpenStudy (anonymous):

Okay, thanks for your effort though.

OpenStudy (anonymous):

np if u ever need help and i know how to do it i will help u

OpenStudy (anonymous):

@iGreen he might can help tho

OpenStudy (anonymous):

@Mertsj @phi Could you see if I got this correct? 1. The main tank has a radius of 70 feet. What is the volume of the quarter-sphered sized tank? Round your answer to the nearest whole number. You must explain your answer using words, and you must show all work and calculations to receive credit. In the description, we are told that we need to find the volume of a quarter-sphered tank, so in order to do that we need to use this formula (Volume of a sphere): V = 4/3 (3.14) (r)3. Now that we know what formula to use, lets apply this to our problem. V = 4/3 (3.14) (r)3. V = 4/3 (3.14) (70)3. V = 4/3 (3.14) (343000). V = 4/3 (1077,020) V = 1,436,026.67 In the description, we are told to find the volume of a “quarter sized sphere”. We have just found the volume for the whole sphere, so in order to find the volume for a quarter of it we need to divide our total volume by four. 1,436,026.67 / 4 = 359,006.6675 Now we round our answer to the nearest whole number and get a grand total of 359,007. Now we know that the volume of the quarter sized sphere tank is 359,007. Lets move on to the calculation!

OpenStudy (igreen):

Yep, looks good to me.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

don't know that I agree with your reading of the question what is a 1/4 sized tank? if the radius is 70 is a 1/4 sized not 70/4 then find the volume